# Evolution of binaries containing a hot subdwarf and a white dwarf to double white dwarfs, and double detonation supernovae with hypervelocity runaway stars

### v1

@@ 4.1 Effect of the donor mass on the mass transfer rate
Figure 8: Critical properties of systems that undergo He ignition in binaries with an accretor of an initial mass of $0.7\,\mathrm{M_{\odot}}$. The figure shows (a) the density of the He layer at the location of ignition, (b) the total mass of He accumulated on the WD, (c) the time-averaged accretion rate from the time when the accretion rate was greater than $\mathrm{10^{-10}}\,\mathrm{M_{\odot}\,yr^{-1}}$ until the He ignition and (d) the peak accretion rate at which the WD gains mass. Systems classified as double detonation supernovae experience steady accretion with lower average and peak rates, leading to the accumulation of more He and ignition at a denser ($>10^{6}\,\mathrm{g\,\text{cm}^{-3}}$) layer.
@@
Fig. 7 illustrates the HR diagram for the evolution of hot subdwarf stars in isolation. Low-mass hot subdwarf stars ($<0.6\,\mathrm{M_{\odot}}$) begin by burning He in the core, enter a thermally pulsing phase once core He is exhausted, and eventually evolve into WDs with thick He shells. In our MESA models, these stars experience periodic thermal pulses because the He shell burning is unstable. The pulses cause the star to undergo significant expansions and contractions, leading to variability in the luminosity. Conversely, high-mass hot subdwarf stars ($>0.6\,\mathrm{M_{\odot}}$) start with core He burning and transition to stable He shell burning upon exhausting core He. After the shell-burning phase, these stars also become WDs with thick He shells.
@@
For accretor and donor masses $<0.75\,\mathrm{M_{\odot}}$, we observe an average accretion rate of approximately $10^{-8}\,\mathrm{M_{\odot}\,yr^{-1}}$ during the core He burning phase of the donor (see panel (c) of Fig. 8). During the He shell burning, thermally pulsing phase, and proto-WD (where donor’s luminosity comes from residual heat rather than fusion) phase, the peak mass transfer rate ranges from about $10^{-7}\,\mathrm{M_{\odot}\,yr^{-1}}$ to $10^{-6}\,\mathrm{M_{\odot}\,yr^{-1}}$ (see panel (d) of Fig. 8). If mass transfer starts during the core He burning phase, the accretion rates are generally slow enough to allow accumulation of large He shells that result in He ignition in the deeper, denser ($\mathrm{\rho>10^{6}\,g\,cm^{-3}}$) (see panel (a) of Fig. 8) layers of the accretor, resulting in a double detonation supernova. Conversely, if mass transfer occurs during the He shell burning, thermally pulsing phase, or proto-WD phase of the donor, He ignition occurs in the less dense ($\mathrm{\rho<10^{6}\,g\,cm^{-3}}$) (see panel (a) of Fig. 8) layers of the accretor, leading to a He nova. For donor masses $>0.75\,\mathrm{M_{\odot}}$, where there is the possibility of the donor mass being greater than that of the accretor in low mass accretors (for example, a $0.7\,\mathrm{M_{\odot}}$ WD and a $0.75\,\mathrm{M_{\odot}}$ WD), the mass transfer rate depends on the mass of the accretor (for more details, see Sec. 4.2).
In our simulations, mass transfer ceases only when the donor exhausts its He or when we stop the simulation when it reaches one of the specified outcomes. In the following subsection, we use systems containing 0.7 WD as examples and explain them in detail.
4.1.1 Explanation to evolutionary fates with a $0.7\,\mathrm{M}_{\odot}$ white dwarf
@@
Panel (a) of Fig. 6 displays outcomes from a grid of binary systems featuring a $0.7\,\mathrm{M_{\odot}}$ accretor and various donors at different initial masses and binary orbital periods. Fig. 8 displays the critical properties of the systems that undergo He ignition in binaries with an accretor of an initial mass of $0.7\,\mathrm{M_{\odot}}$. The figure illustrates (a) the density of the He layer at the location of ignition, (b) the total mass of He accumulated on the WD, (c) the time-averaged accretion rate for times when the instantaneous accretion rate was greater than $\mathrm{10^{-10}}\,\mathrm{M_{\odot}\,yr^{-1}}$, and (d) the peak accretion rate at which the WD gains mass.
@@
First, we describe systems containing low-mass hot subdwarf donors (ranging from $0.33\,\mathrm{M_{\odot}}$ to $0.40\,\mathrm{M_{\odot}}$). For these systems, He ignition occurs in WDs when the initial orbital periods are less than 3 hours. In contrast, no He ignition occurs for initial orbital periods longer than 3 hours. Of the binaries where the accretor undergoes He ignition, the system comes into contact and starts transferring mass during the core He burning phase of the donor and continues as the donor evolves into a WD. Throughout the core He burning phase, they accrete at the rate of $\sim 10^{-8}\,\mathrm{M_{\odot}\,\text{yr}^{-1}}$. When the donor transitions into a proto-WD, the mass transfer rates peak at $\sim 10^{-6}\,\mathrm{M_{\odot}\,\text{yr}^{-1}}$, resulting in He ignition in less dense ($<10^{6}\,\mathrm{g\,\text{cm}^{-3}}$) layers of the WD. This ignition occurs at densities less than the critical density (see Sec. 2.3), classifying all of them as He novae. Systems with low-mass donors ($0.33\,\mathrm{M_{\odot}}$ to $0.40\,\mathrm{M_{\odot}}$) and initial orbital periods longer than $\sim 3$ hours never interact before the hot subdwarf becomes a WD. The initial orbital periods of the double white dwarf systems range from 3 to 7.2 hours. These systems are expected to merge due to gravitational wave radiation in a time ranging from 81 million years to 7.9 billion years. The time until a binary system merges due to gravitational wave radiation (Peters, 1964) is calculated by
@@
where $c$ is the speed of light in a vacuum, and $a$ denotes the semi-major axis of the orbit.
@@
A $0.8\,\mathrm{M_{\odot}}$ donor whose mass is higher than the accretor ($q>1;q=\mathrm{M_{a}}/\mathrm{M_{d}}$) transfers mass at a high average accretion rate ($\sim 10^{-7}\,\mathrm{M_{\odot}\,\text{yr}^{-1}}$) at all initial orbital periods within our range. This high accretion rate ignites the less dense layers of the WD, but the rates are not high enough for the steady burning of He. We therefore classify this system as a He nova. Our initial orbital period range is insufficient for these systems to evolve in isolation to become double white dwarfs.
@@
[…]
@@ 4.2 Effect of the initial accretor mass on the accretion outcome
The initial accretor mass can affect the fate of the binary for two reasons. Firstly, massive WDs have a denser He shell compared to low-mass WDs. For instance, the base of the He shell prior to accretion in our $1.0\,\,\mathrm{M_{\odot}}$ WD is denser ($\rho\sim 2\times 10^{5}\,\mathrm{g\,cm^{-3}}$) compared to that in a $0.7\,\mathrm{M_{\odot}}$ WD ($\rho\sim 0.8\times 10^{5}\,\mathrm{g\,cm^{-3}}$). He ignition is highly dependent on both temperature and density. As a result, the mass of the He required to ignite the $1.0\,\mathrm{M_{\odot}}$ accretor is significantly lower than that required for the $0.7\,\mathrm{M_{\odot}}$ accretor. As a result, for a given accretion rate, the time needed to ignite He is the $1.0\,\mathrm{M_{\odot}}$ accretor is shorter compared to the $0.7\,\mathrm{M_{\odot}}$ accretor. This effect is explained below by comparing binaries with two different accretors $0.7\,\mathrm{M_{\odot}}$ and $1.0\,\mathrm{M_{\odot}}$ but with the same donor mass $0.4\,\mathrm{M_{\odot}}$ and same initial orbital period of 36 minutes.
@@
Secondly, the mass ratio has a direct effect on the mass transfer rate. The mass ratio affects the response of the Roche radius to the mass transfer and hence the mass accretion rate. Bauer & Kupfer (2021) studies the phases of mass transfer in hot subdwarf + WD binaries. In particular, they present an analytic expression to estimate the Roche radius response to the mass transfer as $({\mathrm{d}\log R_{\text{RL}}}/{\mathrm{d}\log M_{d}})\approx-\frac{5}{3}+2.1q$. Let us compare binaries with two different initial accretor masses, $0.7\,\mathrm{M_{\odot}}$ and $1.0\,\mathrm{M_{\odot}}$, but with the same donor of $0.8\,\mathrm{M_{\odot}}$ and same initial period of 36 minutes. For an accretor of $0.7\,\mathrm{M_{\odot}}$ and a donor of $0.8\,\mathrm{M_{\odot}}$, the mass ratio $q$ is greater than 1, but for a $1.0\,\mathrm{M_{\odot}}$ accretor, $q<1$. For a larger mass ratio as in $0.7\,\mathrm{M_{\odot}}$ accretor and a $0.8\,\mathrm{M_{\odot}}$ donor, the Roche lobe contracts as the donor loses mass, resulting in an increased mass transfer rate. Whereas for a $1\,\mathrm{M_{\odot}}$ accretor, the Roche radius of the donor increases in response to mass loss, leading to a relatively smaller accretion rate, leading to a double detonation supernova.
@@ 4.3 Outliers
In this section, we present two interesting outlier systems that deviate from the overall trends of the grid in Fig. 6. The detailed evolution diagram can be found in the appendix A.
@@ 5.1 He shell Masses at ignition
Figure 10: Orbital velocities of the donor at the time of the explosion of the WD. The blue line and red line indicate inferred ejection velocities of the observed He-rich stars US708 (Neunteufel 2020) and J2050 (Ziegerer et al. 2017) respectively. The dark and light green shaded regions represent the $1\,\sigma$ and $2\,\sigma$ uncertainties to J2050. Low-mass donors with high-mass accretors result in the highest possible orbital velocities. The outliers with higher velocities for the same initial donor mass indicate the donor is a proto-WD at the time of He ignition.
He-accreting WDs are one of the progenitors of double detonation supernovae. Sub-Chandrasekhar explosions can occur when explosive burning in the He shell of the WDs triggers a second detonation within the WD, resulting in a thermonuclear supernova. However, the He mass required for the first detonation is uncertain. Previous studies (Woosley & Kasen, 2011; Sim et al., 2012) suggest that in a single degenerate scenario, a He shell of 0.1 - 0.2 $\,\mathrm{M_{\odot}}$ is required to trigger the first detonation. But such a massive He shell would lead to the production of titanium, chromium, and nickel, leading to deviation from normal supernova Ia spectra (Kromer et al., 2010; Sim et al., 2012; Shen et al., 2021; Collins et al., 2022). Here, we present the He shell masses for the systems that undergo double detonation supernovae.
@@ 5.2 Runaway velocities
A He-rich star, US 708, was originally observed by Usher et al. (1982) and later observed by SDSS. Hirsch et al. (2005) measured the velocity of the star and suggested that its high velocity might result from interactions in the Galactic center. Justham et al. (2009) proposed that this velocity could be attributed to the orbital velocity of a hot subdwarf star after its WD companion underwent a Type Ia supernova explosion. This hypothesis was supported by Geier et al. (2015), and Brown et al. (2015), who demonstrated that the trajectory of the hot subdwarf star cannot be traced back to the Galactic center. Given the increasing number of hypervelocity runaway stars discovered by Gaia and the upcoming data releases, we anticipate more detections of these stars.
Since our grid includes models that might lead to WD explosions, we present here the orbital velocities of these donor stars at the time of the explosion of the WD. The orbital velocities range from $\sim 453\,\mathrm{km\,s^{-1}}$ to $\sim 1018\,\mathrm{km\,s^{-1}}$. Fig. 10 shows the orbital velocity of these donor stars as a function of hot subdwarf mass. The colors represent the current white dwarf mass at the time of the explosion. Neunteufel (2020) estimated the ejection velocity of US 708 after taking into account the updated proper motion measurements from Gaia Collaboration et al. (2018). With the inferred ejection velocity to be 894 $\mathrm{km\,s^{-1}}$, and assuming a Chandrasekhar mass supernova, they estimate the mass of US 708 to be in the range $0.34\,\mathrm{M_{\odot}}<M_{\mathrm{US708}}<0.37\,\mathrm{M_{\odot}}$ (see the blue line in Fig. 10). In Fig. 10 the red line and green shaded regions show the estimated ejection velocities of J2050 (a spectroscopic twin of US 708; Ziegerer et al. 2017), which are $385\pm 78\,\mathrm{km\,s^{-1}}$. From our models, there is one surviving donor with a mass of $0.63\,\mathrm{M_{\odot}}$ that lies within their $1\,\sigma$ range, and 69 models within the mass range $0.5\,\mathrm{M_{\odot}}<M_{\mathrm{d}}<0.73\,\mathrm{M_{\odot}}$ fall within their $2\,\sigma$ range.
The He star is expected to receive kick velocities of 10-100 $\mathrm{km\,s^{-1}}$ from the ejecta (Liu et al., 2023). However, this is much lower compared than the pre-explosion orbital velocity of the star. Hence, the runway velocities are predominantly determined by the pre-explosion orbital velocities (Liu et al., 2023; Braudo & Soker, 2024). A proto-WD can also be ejected from these binaries if the mass transfer phase is prolonged or begins at a later phase, allowing the donor sufficient time to advance in its evolution toward a WD phase. In Fig. 10, the outliers that have higher velocities than hot subdwarfs of the same mass represent these proto-WDs.

### v2

@@ 4.1 Effect of the donor mass on the mass transfer rate
Figure 8: Critical properties of systems that undergo He ignition in binaries with an accretor of an initial mass of $0.7\,\mathrm{M_{\odot}}$. The figure shows (a) the density of the He layer at the location of ignition, (b) the total mass of He accumulated on the WD, (c) the time-averaged accretion rate from the time when the accretion rate was greater than $\mathrm{10^{-10}}\,\mathrm{M_{\odot}\,yr^{-1}}$ until the He ignition and (d) the peak accretion rate at which the WD gains mass. Systems classified as double detonation supernovae experience steady accretion with lower average and peak rates, leading to the accumulation of more He and ignition at a denser ($10^{6}\,\mathrm{g\,\text{cm}^{-3}}$) layer.
@@
Fig. 7 illustrates the HR diagram for the evolution of hot subdwarf stars in isolation. Low-mass hot subdwarf stars ($0.6\,\mathrm{M_{\odot}}$) begin by burning He in the core, enter a thermally pulsing phase once core He is exhausted, and eventually evolve into WDs with thick He shells. In our MESA models, these stars experience periodic thermal pulses because the He shell burning is unstable. The pulses cause the star to undergo significant expansions and contractions, leading to variability in the luminosity. Conversely, high-mass hot subdwarf stars ($0.6\,\mathrm{M_{\odot}}$) start with core He burning and transition to stable He shell burning upon exhausting core He. After the shell-burning phase, these stars also become WDs with thick He shells.
@@
For accretor and donor masses $0.75\,\mathrm{M_{\odot}}$, we observe an average accretion rate of approximately $10^{-8}\,\mathrm{M_{\odot}\,yr^{-1}}$ during the core He burning phase of the donor (see panel (c) of Fig. 8). During the He shell burning, thermally pulsing phase, and proto-WD (where donor’s luminosity comes from residual heat rather than fusion) phase, the peak mass transfer rate ranges from about $10^{-7}\,\mathrm{M_{\odot}\,yr^{-1}}$ to $10^{-6}\,\mathrm{M_{\odot}\,yr^{-1}}$ (see panel (d) of Fig. 8). If mass transfer starts during the core He burning phase, the accretion rates are generally slow enough to allow accumulation of large He shells that result in He ignition in the deeper, denser ($\mathrm{\rho>10^{6}\,g\,cm^{-3}}$) (see panel (a) of Fig. 8) layers of the accretor, resulting in a double detonation supernova. Conversely, if mass transfer occurs during the He shell burning, thermally pulsing phase, or proto-WD phase of the donor, He ignition occurs in the less dense ($\mathrm{\rho<10^{6}\,g\,cm^{-3}}$) (see panel (a) of Fig. 8) layers of the accretor, leading to a He nova. For donor masses $0.75\,\mathrm{M_{\odot}}$, where there is the possibility of the donor mass being greater than that of the accretor in low mass accretors (for example, a $0.7\,\mathrm{M_{\odot}}$ WD and a $0.75\,\mathrm{M_{\odot}}$ WD), the mass transfer rate depends on the mass of the accretor (for more details, see Sect. 4.2).In our simulations, mass transfer ceases only when the donor exhausts its He or when we stop the simulation when it reaches one of the specified outcomes.
@@
We use systems containing 0.7 WD as examples and explain them in detail. Panel (a) of Fig. 6 displays outcomes from a grid of binary systems featuring a $0.7\,\mathrm{M_{\odot}}$ accretor and various donors at different initial masses and binary orbital periods. Fig. 8 displays the critical properties of the systems that undergo He ignition in binaries with an accretor of an initial mass of $0.7\,\mathrm{M_{\odot}}$. The figure illustrates (a) the density of the He layer at the location of ignition, (b) the total mass of He accumulated on the WD, (c) the time-averaged accretion rate for times when the instantaneous accretion rate was greater than $\mathrm{10^{-10}}\,\mathrm{M_{\odot}\,yr^{-1}}$, and (d) the peak accretion rate at which the WD gains mass.
@@
First, we describe systems containing low-mass hot subdwarf donors (ranging from $0.33\,\mathrm{M_{\odot}}$ to $0.40\,\mathrm{M_{\odot}}$). For these systems, He ignition occurs in WDs when the initial orbital periods are less than 3 hours. In contrast, no He ignition occurs for initial orbital periods longer than 3 hours. Of the binaries where the accretor undergoes He ignition, the system comes into contact and starts transferring mass during the core He burning phase of the donor and continues as the donor evolves into a WD. Throughout the core He burning phase, they accrete at the rate of $\sim 10^{-8}\,\mathrm{M_{\odot}\,\text{yr}^{-1}}$. When the donor transitions into a proto-WD, the mass transfer rates peak at $\sim 10^{-6}\,\mathrm{M_{\odot}\,\text{yr}^{-1}}$, resulting in He ignition in less dense ($10^{6}\,\mathrm{g\,\text{cm}^{-3}}$) layers of the WD. This ignition occurs at densities less than the critical density (see Sect. 2.1), classifying all of them as He novae. Systems with low-mass donors ($0.33\,\mathrm{M_{\odot}}$ to $0.40\,\mathrm{M_{\odot}}$) and initial orbital periods longer than $\sim 3$ hours never interact before the hot subdwarf becomes a WD. The initial orbital periods of the double white dwarf systems range from 3 to 7.2 hours. These systems are expected to merge due to gravitational wave radiation in a time ranging from 81 million years to 7.9 billion years. The time until a binary system merges due to gravitational wave radiation (1964PhRv..136.1224P) is calculated by
@@
where and $a$ denotes the semi-major axis of the orbit.
@@
A $0.8\,\mathrm{M_{\odot}}$ donor whose mass is higher than the accretor ($q>1;q=\mathrm{M_{a}}/\mathrm{M_{d}}$) transfers mass at a high average accretion rate ($\sim 10^{-7}\,\mathrm{M_{\odot}\,\text{yr}^{-1}}$) at all initial orbital periods within our range. This high accretion rate ignites the less dense outer layers of the white dwarf. However, the rates are not sufficient to sustain steady helium burning and subsequent C/O core growth. We therefore classify this system as a He nova. Our initial orbital period range is insufficient for these systems to evolve in isolation to become double white dwarfs.
@@
Figure 9: He masses (panel a) and mass coordinates (panel b) for systems that undergo double detonation supernovae and He novae. Panel (a) shows the mass of the He shell at the time of ignition versus the total mass of the WD for systems undergoing double detonation supernovae and He novae. The minimum He mass required for He ignition in our simulations, at high densities greater than the assumed critical density is $\sim 0.05\,\mathrm{M_{\odot}}$. Panel (b): Red stars with black edges display the mass coordinate of the ignition point versus the mass coordinate of the C/O–He boundary, where the C/O-He boundary is defined as the layer containing less than $10\%$ He. The red dashed line represents the locus where the mass coordinate of the C/O-He boundary equals that of the ignition point.
@@ 4.2 Effect of the initial accretor mass on the accretion outcome
The initial accretor mass can affect the fate of the binary for two reasons. Firstly, massive WDs have a denser He shell compared to low-mass WDs. For instance, the base of the He shell prior to accretion in our $1.0\,\,\mathrm{M_{\odot}}$ WD is denser ($\rho\sim 2\times 10^{5}\,\mathrm{g\,cm^{-3}}$) compared to that in a $0.7\,\mathrm{M_{\odot}}$ WD ($\rho\sim 0.8\times 10^{5}\,\mathrm{g\,cm^{-3}}$). He ignition is highly dependent on both temperature and density. As a result, the mass of the He required to ignite the $1.0\,\mathrm{M_{\odot}}$ accretor is significantly lower than that required for the $0.7\,\mathrm{M_{\odot}}$ accretor. As a result, for a given accretion rate, the time needed to ignite He in the $1.0\,\mathrm{M_{\odot}}$ accretor is shorter compared to the $0.7\,\mathrm{M_{\odot}}$ accretor. This effect is explained below by comparing binaries with two different accretors $0.7\,\mathrm{M_{\odot}}$ and $1.0\,\mathrm{M_{\odot}}$ but with the same donor mass $0.4\,\mathrm{M_{\odot}}$ and same initial orbital period of 36 minutes.
@@
Secondly, the mass ratio has a direct effect on the mass transfer rate. The mass ratio affects the response of the Roche radius to the mass transfer and hence the mass accretion rate. 2021ApJ...922..245B studied the phases of mass transfer in hot subdwarf + WD binaries. In particular, they present an analytic expression to estimate the Roche radius response to the mass transfer as $({\mathrm{d}\log R_{\text{RL}}}/{\mathrm{d}\log M_{d}})\approx-\frac{5}{3}+2.1q$. Let us compare binaries with two different initial accretor masses, $0.7\,\mathrm{M_{\odot}}$ and $1.0\,\mathrm{M_{\odot}}$, but with the same donor of $0.8\,\mathrm{M_{\odot}}$ and same initial period of 36 minutes. For an accretor of $0.7\,\mathrm{M_{\odot}}$ and a donor of $0.8\,\mathrm{M_{\odot}}$, the mass ratio $q$ is greater than 1, but for a $1.0\,\mathrm{M_{\odot}}$ accretor, $q<1$. For a larger mass ratio as in $0.7\,\mathrm{M_{\odot}}$ accretor and a $0.8\,\mathrm{M_{\odot}}$ donor, the Roche lobe contracts as the donor loses mass, resulting in an increased mass transfer rate. Whereas for a $1\,\mathrm{M_{\odot}}$ accretor, the Roche radius of the donor increases in response to mass loss, leading to a relatively smaller accretion rate, leading to a double detonation supernova.
@@ 4.3 Outliers
In this section, we present two interesting outlier systems that deviate from the overall trends of the grid in Fig. 6. The detailed evolution diagram can be found in appendix A.
@@ 5.1 He shell masses at ignition
Figure 10: Orbital velocities of the donor at the time of the explosion of the WD. The blue line and red line indicate inferred ejection velocities of the observed He-rich stars US708 (2020A&A...641A..52N) and J2050 (2017A&A...601A..58Z) respectively. The dark and light green shaded regions represent the $1\,\sigma$ and $2\,\sigma$ uncertainties to J2050. Low-mass donors with high-mass accretors result in the highest possible orbital velocities. The outliers with higher velocities for the same initial donor mass indicate the donor is a proto-WD at the time of He ignition.
He-accreting WDs are one of the progenitors of double detonation supernovae. Sub-Chandrasekhar explosions can occur when explosive burning in the He shell of the WDs triggers a second detonation within the WD, resulting in a thermonuclear supernova. However, the He mass required for the first detonation is uncertain. Previous studies (2011ApJ...734...38W; 2012MNRAS.420.3003S) suggest that in a single degenerate scenario, a He shell of 0.1 - 0.2 $\mathrm{M_{\odot}}$ is required to trigger the first detonation. But such a massive He shell would lead to the production of titanium, chromium, and nickel, leading to deviation from normal supernova Ia spectra (2010ApJ...719.1067K; 2012MNRAS.420.3003S; 2021ApJ...922...68S; 2022MNRAS.517.5289C). Here, we present the He shell masses for the systems that undergo double detonation supernovae.
@@ 5.2 Runaway velocities
A He-rich star, US 708, was originally observed by 1982ApJS...48...51U and later observed by SDSS. 2005A&A... [^4] L..61H measured the velocity of the star and suggested that its high velocity might result from interactions in the Galactic center. 2009A&A...493.1081J proposed that this velocity could be attributed to the orbital velocity of a hot subdwarf star after its WD companion underwent a Type Ia supernova explosion. This hypothesis was supported by 2015Sci...347.1126G, and 2015ApJ...804...49B, who demonstrated that the trajectory of the hot subdwarf star cannot be traced back to the Galactic center. Given the increasing number of hypervelocity runaway stars discovered by Gaia and the upcoming data releases, we anticipate more detections of these stars.
Since our grid includes models that might lead to WD explosions, we present here the orbital velocities of these donor stars at the time of the explosion of the WD. The orbital velocities range from $\sim 453\,\mathrm{km\,s^{-1}}$ to $\sim 1018\,\mathrm{km\,s^{-1}}$. Fig. 10 shows the orbital velocity of these donor stars as a function of hot subdwarf mass. The colors represent the current white dwarf mass at the time of the explosion. We note that the mass of the runaway stars may vary depending on the supernova ejecta-donor interaction. 2020A&A...641A..52N estimated the ejection velocity of US 708 after taking into account the updated proper motion measurements from 2018A&A...616A...1G. With the inferred ejection velocity to be 894 $\mathrm{km\,s^{-1}}$, and assuming a Chandrasekhar mass supernova, they estimate the mass of US 708 to be in the range $0.34\,\mathrm{M_{\odot}}<M_{\mathrm{US708}}<0.37\,\mathrm{M_{\odot}}$ (see the blue line in Fig. 10). In Fig. 10 the red line and green shaded regions show the estimated ejection velocities of J2050 (a spectroscopic twin of US 708; 2017A&A...601A..58Z), which are $385\pm 78\,\mathrm{km\,s^{-1}}$. From our models, there is one surviving donor with a mass of $0.63\,\mathrm{M_{\odot}}$ that lies within their $1\,\sigma$ range, and 69 models within the mass range $0.5\,\mathrm{M_{\odot}}<M_{\mathrm{d}}<0.73\,\mathrm{M_{\odot}}$ fall within their $2\,\sigma$ range.
The He star is expected to receive kick velocities of 10-100 $\mathrm{km\,s^{-1}}$ from the ejecta (2023RAA....23h2001L). However, this is much lower compared than the pre-explosion orbital velocity of the star. Hence, the runway velocities are predominantly determined by the pre-explosion orbital velocities (2023RAA....23h2001L; 2024OJAp....7E...7B). A proto-WD can also be ejected from these binaries if the mass transfer phase is prolonged or begins at a later phase, allowing the donor sufficient time to advance in its evolution toward a WD phase. In Fig. 10, the outliers that have higher velocities than hot subdwarfs of the same mass represent these proto-WDs. Some of the faster proto-white dwarfs could potentially account for hypervelocity stars with ejection velocities $\lesssim 1000\ \mathrm{km/s}$, such as those reported by 2023OJAp....6E..28E, although confirming this would require modeling the impact of supernova ejecta on these stars.